I am beyond confused on how to solve system of equations.. The book example is not clear and there Is Only one example! The question is :



Solve the following systems of equations. If the system cannot be solved, state that it cannot be solved and explain why.

3x+y=7
3x+y=10

Answer by josgarithmetic(39617)

Same system, x must be equal in both equations at the same time; y must be equal in both equations at the same time.
One equation says 3x+y=7 and the other equation says 3x+y=10. There is a transitive property of equality which says that a=b and b=c means that a=c. This property must seem intuitively correct and therefore perfectly understood.
Using this transitive property for your given system, 7 is equal to the binomial expression 3x+y; and 10 is equal to the binomial expression 3x+y. If 7=3x+y and if 3x+y=10, then according to transitive property, 7=10. You see something wrong now? At least you should understand that these two equations cannot be in the same system of equations.